# Tree Proof-of-Position Algorithms

### v1

@@ Abstract
We present a novel class of proof-of-position algorithms: Tree-Proof-of-Position (T-PoP). This algorithm is decentralised, collaborative and can be computed in a privacy preserving manner, such that agents do not need to reveal their position publicly. We make no assumptions of honest behaviour in the system, and consider varying ways in which agents may misbehave. Our algorithm is therefore resilient to highly adversarial scenarios. This makes it suitable for a wide class of applications, namely those in which trust in a centralised infrastructure may not be assumed, or high security risk scenarios. Our algorithm has a worst case quadratic runtime, making it suitable for hardware constrained IoT applications. We also provide a mathematical model that summarises T-PoP’s performance for varying operating conditions. We then simulate T-PoP’s behaviour with a large number of agent-based simulations, which are in complete agreement with our mathematical model, thus demonstrating its validity. T-PoP can achieve high levels of reliability and security by tuning its operating conditions, both in high and low density environments. Finally, we also present a mathematical model to probabilistically detect platooning attacks.
@@ I Introduction
An emerging problem in many domains concerns the verification that an individual, or an object, is in the location that it claims to be in. Examples of situations where such a need arises range from images in conflict zones to vehicles who need to prove their location in order to access services. Proof-of-position algorithms refer to a suite of algorithmic tools that are designed to enable such objects to prove their location.
@@
Hence we advocate for the need for a proof-of-position algorithm that is decentralised. However, decentralisation alone does not suffice. Information on an individual’s position is sensitive. Revealing it may incur safety risks to the user. Thus, the proof-of-position algorithms must also be privacy preserving. More concisely: one should be able to provide a verifiable proof of their position, without having to publicly reveal their position. In addition, for any solution to be applicable to real-world use cases, it must be designed to withstand adversarial environments and account for individuals that lie and misbehave. In the context of proof-of-position, there are many ways an agent could lie: they may lie about their own position, but also about the position of others. With these considerations in mind, our objective in this paper is to provide a proof-of-position algorithm that is decentralised; private; and robust to adversarial attacks.
@@ II Related Work
A number of works have appeared in the literature addressing the problem of proving one’s position; see for example [3], [4], [5], [6], [7]. These proposals are unsuitable for high risk scenarios we consider because they introduce trust assumptions that may not hold in adversarial environments. They also de facto re-introduce a degree of centralisation in their systems. We proceed by outlining some of this prior work.
An early example of a decentralized proof-of-location scheme, known as APPLAUS, was introduced in [8]. APPLAUS addressed collusion attacks through graph clustering and computing a ‘betweeness’ metric to evaluate node trustworthiness. In [9], nodes weakly connected in the graph are considered less reliable. They propose a time-decayed weight function and determine node trustworthiness based on the ratio of approvals to neighbours. This latter contribution serves as a starting point for our work. Whilst offering a number of valuable contributions, this work relies on a central server that stores information on the number of agents at particular time and location to detect fraudulent proof generation. This information can be used to estimate whether a prover lies about not finding enough peers or always finding the same peer based on some statistical techniques. Conversely, the security and integrity of our algorithm does not rely assuming access to a trusted central server that stores said information.
@@
The authors in [2] propose a proof-of-position protocol using blockchain agnostic smart contracts. A number of proposals have emerged to design proof-of-position schemes leveraging blockchain technologies, see: [7], [6], [13], [14]. Blockchain is known to centralise over time [15], and is highly resource intensive [16]. Our protocol does not rely on blockchain, thus avoiding the drawbacks inherent to its design.
Contributions with respect to prior-art: In this paper we present a decentralised proof-of-position algorithm, T-PoP. It is decentralised because it does not rely on a trusted central server or authority to ensure the position proof is valid. It is privacy preserving, because the position proof may be computed without revealing the position of an agent publicly, and other participants may verify that the position proof is indeed valid. Indeed, we provide a proof of concept using Zama’s fully homomorphic encryption library [17], where agents provide their encrypted position, and run T-PoP without having to decrypt it. It is available on our GitHub repository. Our algorithm is also robust to highly adversarial scenarios: we consider that agents may not only lie about their own position, but also about the position of other agents. We take this consideration into account to provide a probabilistic guarantee of an agent’s position. We provide a mathematical model that summarises the behaviour of our algorithm, and validate that the model does indeed represent our algorithm through the use of agent-based simulations. This model provides designers with security and reliability guarantees of T-PoP’s performance, given an expected distribution of malicious agents in the network. Thus, designers may use this model as opposed to conducting resource intensive agent based simulations to determine what operating conditions of our algorithm best suit their needs. Our algorithm runs in quadratic time and thus can be implemented in a wide range of IoT applications. T-PoP is a solution to prove one’s location in a highly adversarial and mutually distrusting environment, where a central authority and trusted infrastructure cannot be relied upon. We demonstrate it performs well in both high and low density applications, given that its operating conditions may be tuned to provide increased security or reliability.
@@ III Components
Agents: Any participant in the proof-of-position protocol is considered an agent, $a_{i}$, where $a_{i}\in A$, $|A|=N$ and $N$ is the total number of agents operating in the protocol. All agents are in a given state, that determines if they are honest, and coerced, or not. For brevity, we omit the agent’s index where unnecessary.
@@
Position: All agents have a real position, $p$, and a claimed position, $\hat{p}$. Each agent can lie about their claimed position by committing to $\hat{p}$ such that $\hat{p}\neq p$.
$\hat{p}$$\in\{p,p^{\prime}\}$ (1a)
$\hat{p}=p$ $\vee\hat{p}=p^{\prime}$ (1b)
$p^{\prime}$ $\neq p$ (1c)
In other words, the claimed position may or may not be equal to the real position. We say that $\{\hat{p},p,p^{\prime}\}\in E$.
@@
Note: this field of view is not restrictive of physical obstacles blocking immediate range of sight, rather, it is intended as a model for connectivity range.
@@ V Implementation details of T-PoP
T-PoP’s practical implementations will vary depending on the application. It may be used in by vehicles to prove their position within a city, or it could be implemented in a smartphone, operated by individuals wishing to prove their location. It is also worth noting that T-PoP is not a highly resource intensive algorithm to execute, and thus the hardware constraints are not a critical concern. It is comprised of two algorithms: Verification and Checks. These are further defined in 3 and 2 accordingly. Verification is a (reverse) breadth-first-search algorithm, meaning its runtime in the worst case scenario is $O(|V|+|E|)$, where $V$ is the number of nodes and $E$ the number of edges in the tree data structure [22]. The Verification algorithm calls the Checks algorithm at each node, and the Checks algorithm has a time complexity of $O(|V|)$. So T-PoP’s time complexity is $O(|V|^{2}+|E|)$.
@@ VII-A Agent States
. In practical terms: an honest agent will always claim to be in their real position. If $P(\alpha_{i}=\overline{h})$, finding $P(\gamma_{i}=t)$ is slightly more convoluted. This is because when a dishonest agent is sampled in a tree, one cannot, in theory, distinguish if its claimed position is its real position or not. Indeed, that is the purpose of T-PoP: to determine whether an agent is claiming to be in their real position.
@@ VII-B2 Picking a child, given a parent
$x(\textbf{c})=s_{j}=\{\alpha_{\textbf{c}},\beta_{\textbf{c}},\gamma_{\textbf{c}}\}$
$x(\textbf{p})=s_{i}=\{\alpha_{\textbf{p}},\beta_{\textbf{p}},\gamma_{\textbf{p}}\}$
To compute $m_{i,j}$ it suffices to evaluate:
@@
. Note that only the coercion attribute of the parent affects the result of equations 18. We include all three attributes to ease exposition, to show all possible state combinations.
@@ VII-D Condition 3: all the nodes in the tree are unique
  • A tree with height $h=1$ and branching factor $w=6$.
  • A tree with height $h=2$ and branching factor $w=2$.
We denote this probability as $P(h,w)$. In the remainder of this section we make the assumption that each agent selects witnesses among its neighbours without replacement. Although it may be an unrealistic assumption (as in practice each prover would select distinct witnesses and they would not deliberately try to hinder their own proof), it allows us to provide a lower bound on the probability of selecting a tree made of unique nodes.
@@
For the case of $h=1$ and $w=6$, $P(h,w)$ is straightforward to compute and it is equal to $P(1,6)=\kappa_{0,6}$.
For the tree of $h=2$ and $w=2$, finding the probability that all nodes are unique is more complex. Let us consider a tree started by a prover $a_{1}$ at the root, tree height of $h=2$, and branching factor of $w=2$, i.e., each parent choosing two of its neighbours at random (with replacement) as its children (i.e., the nodes at level 1 will not choose themselves).
@@
Finally, inserting equation (30) into equation (27), we have the theoretical approximation for the probability of the tree ($h=w=2$) being comprised of strictly distinct agents, i.e., $P(2,2)=P(L_{0},L_{1},L_{2})=\kappa_{0,2}\cdot P(L2|L_{0},L_{1})$. This value of $P(L_{0},L_{1},L_{2})$, given a constant range of sight $r$, is only dependent on the density of agents $N/l^{2}$ and can be calculated numerically.
@@
As a final remark we note that while it may be possible to obtain a a recursive formula to compute the probability of finding distinct agents in a tree with depth $h$ and branching factor $w$, this might prove very complicated and it is quite straightforward to approximate $P(h,w)$, for a given $\langle k\rangle$ and a given $r$, by making use of a large number of Monte-Carlo simulations.
@@ IX Agent Based Simulations
We construct an agent-based simulator, coded in Python, which is available in our GitHub Repository. In it, we create an environment of a given size, and instantiate a fixed number of agents with a fixed value for their range of sight. Agents are assigned an honesty, $\alpha_{i}$, and a coercion, $\beta_{i}$, attribute independently and at random, following equations 10a and 10b. All agents are given a real position, within the bounds of the defined environment, and dishonest agents also have a claimed position that is different to their real position. These positions are uniformly and randomly distributed across the environment. We first simulate a low-density scenario, as shown in Figures 17, 16, 19 and 18. We initialise an environment of size 10 by 10 arbitrary square units, and instantiate 350 agents, such that each has a range of sight of 1 unit. Thus, agents have an average of 3.5 agents within their field of view. Each agent constructs a tree, and then runs the T-PoP algorithms to obtain a proof-of-position. These simulations are carried out across the ranging values of $p_{h}$ and $p_{c}$, starting from 0 to 100% in intervals of 10%. At each combination of $p_{h}$ and $p_{c}$, 5 Monte Carlo simulations are run for every agent. In total, that is 211,750 T-PoP simulations [^4] There are 121 possible combinations of $p_{h}$ and $p_{c}$ when varying the values at intervals of 10%, 350 agents and 5 simulations per possible combination of $p_{h}$ and $p_{c}$ yields $5*121*350$.. We test the performance of T-PoP for the following operating conditions:
$threshold=100\%,w_{d}=2,h=2$
$threshold=100\%,w_{d}=6,h=1$
$threshold=40\%,w_{d}=2,h=2$
$threshold=40\%,w_{d}=6,h=1$
We measure the recall rate of true positives, true negatives, false positives and false negatives. Each classification instance is described below:
@@
We consider recall over precision, since we are particularly interested in the sensitivity of the algorithm. Namely, it is of utmost importance that the number of False Positives is low, as this is the most dangerous possible outcome.
@@ X Performance of TPoP
From the results obtained, it can be observed that the performance of TPoP does not vary greatly when using a tree of $h=1,w=6$ versus $h=2,w=2$. The prior slightly out-performs the latter in both thresholds tested. This is consistent with the results shown in Figures 10. The main difference in performance is observed when varying the threshold parameter, and when varying the density. Increasing density increases the number of True Positives, given that honest agents will not be classified as dishonest for not having sufficient nodes in their tree. Decreasing the threshold improves the rate of True Positives (Reliability) of T-PoP in very low density scenarios, but this comes at the expense of Security. A lower threshold makes it easier for dishonest agents to submit a valid tree. As such, users may tune the threshold parameters to obtain a desired level of Security or Reliability, depending on the expected density of their system.

### v2

@@ Abstract
We present a novel class of proof-of-position algorithms: Tree-Proof-of-Position (T-PoP). This algorithm is decentralised, collaborative and can be computed in a privacy preserving manner, such that agents do not need to reveal their position publicly. We make no assumptions of honest behaviour in the system, and consider varying ways in which agents may misbehave. Our algorithm is therefore resilient to highly adversarial scenarios. This makes it suitable for a wide class of applications, namely those where trust in a centralised infrastructure may not be assumed, or high security risk scenarios. Our algorithm has a worst case quadratic runtime, making it suitable for hardware constrained IoT applications. We also provide a mathematical model that allows us to analyse T-PoP’s performance for varying operating conditions. We then simulate T-PoP’s behaviour with a large number of agent-based simulations, which are in complete agreement with our mathematical model. T-PoP can achieve high levels of reliability and security by tuning its operating conditions, both in high and low density environments. Finally, we also present a mathematical model to probabilistically detect platooning attacks.
@@ I Introduction
An emerging problem in many domains concerns the verification that an individual, or an object, is in the location that it claims to be in. Examples of situations where such a need arises range from images in conflict zones to vehicles who need to prove their location in order to access services. Proof-of-position algorithms refer to a suite of algorithmic tools that are designed to enable such objects to prove their location. An immediate question that arises is why not use GPS to determine one’s location. GPS signals can be spoofed [1], [2].
@@
Hence we advocate for the need for a proof-of-position algorithm that is decentralised. However, decentralisation alone does not suffice. Information on an individual’s position is sensitive. Revealing it may incur safety risks to the user. Thus, the proof-of-position algorithms must also be privacy preserving. Namely, one should be able to provide a verifiable proof of their position, without having to publicly reveal their position. In addition, for any solution to be applicable to real-world use cases, it must be designed to withstand adversarial environments and account for individuals that lie and misbehave. In the context of proof-of-position, there are many ways an agent could lie: they may lie about their own position, but also about the position of others. With these considerations in mind, our objective in this paper is to provide a proof-of-position algorithm that is decentralised; private; and robust to adversarial attacks.
@@ II Related Work
A number of works have appeared in the literature addressing the problem of proving one’s position; see for example [5, 6, 7, 8, 9]. These proposals are unsuitable for high risk scenarios we consider because they introduce trust assumptions that may not hold in adversarial environments. They also de facto re-introduce a degree of centralisation in their systems. We proceed by outlining some of this prior work.
An early example of a decentralized proof-of-location scheme, known as APPLAUS, was introduced in [10]. APPLAUS addressed collusion attacks through graph clustering and computing a ‘betweeness’ metric to evaluate node trustworthiness. In [11], nodes weakly connected in the graph are considered less reliable. They propose a time-decayed weight function and determine node trustworthiness based on the ratio of approvals to neighbours. This latter contribution serves as a starting point for our work. Whilst offering a number of valuable contributions, their proposal relies on a central server that stores information on the number of agents at particular time and location to detect fraudulent proof generation. This information can be used to estimate whether a prover lies about not finding enough peers or always finding the same peer based on various statistical techniques. Conversely, the security and integrity of our algorithm does not rely assuming access to a trusted central server that stores said information.
@@
The authors in [4] propose a proof-of-position protocol using blockchain agnostic smart contracts. A number of proposals have emerged to design proof-of-position schemes leveraging blockchain technologies, see [9], [8], [15], [16]. Blockchain is known to centralise over time [17], and is highly resource intensive [18]. Our protocol does not rely on blockchain, thus avoiding the drawbacks inherent to its design.
Contributions with respect to prior-art: In this paper we present a decentralised proof-of-position algorithm, T-PoP. It is decentralised because it does not rely on a trusted central server or authority to ensure the position proof is valid. It is privacy preserving, because the position proof may be computed without revealing the position of an agent publicly, and other participants may verify that the position proof is indeed valid. Indeed, we provide a proof of concept using Zama’s fully homomorphic encryption library [19], where agents provide their encrypted position, and run T-PoP without having to decrypt it. It is available on our GitHub repository. Our algorithm is also robust to highly adversarial scenarios: we consider that agents may not only lie about their own position, but also about the position of other agents. We take this consideration into account to provide a probabilistic guarantee of an agent’s position. We mathematically model the behavior of the algorithm, verify the model via agent-based simulations, and predict the algorithm’s behavior in different scenarios analytically by the means of the verified model. This model provides designers with security and reliability guarantees of T-PoP’s performance, given an expected distribution of malicious agents in the network. Thus, designers may use this model as opposed to conducting resource intensive agent based simulations to determine what operating conditions of our algorithm best suit their needs. Our algorithm runs in quadratic time (see section 4) and thus can be implemented in a wide range of IoT applications. T-PoP is a solution to prove one’s location in a highly adversarial and mutually distrusting environment, where a central authority and trusted infrastructure cannot be relied upon. We demonstrate it performs well in both high and low density scenarios, given that its operating conditions may be tuned to provide increased security or reliability.
@@ III Components
Agents: Any participant in the proof-of-position protocol is considered an agent, $a_{i}$, where $a_{i}\in A$, $|A|=N$ and $N$ is the total number of agents operating in the protocol. All agents are in a given state, that determines if they are honest, and coerced, or not. For brevity, we omit the agent’s index where possible.
@@
Position: All agents have a real position, $p$, and a claimed position, $\hat{p}$. Each agent can claim to be in a position different to their real one (i.e.: lie about their position) by committing to $\hat{p}$ such that $\hat{p}\neq p$.
$\hat{p}$ $\neq p\vee\hat{p}=p$ (1a)
In other words, the claimed position may or may not be equal to the real position. We say that $\hat{p},p\in E$.
@@
Note that this field of view is not restrictive of physical obstacles blocking immediate range of sight, rather, it is intended as a model for connectivity range.
@@ V Implementation details of T-PoP
T-PoP’s practical implementations will vary depending on the application. It may be used in by vehicles to prove their position within a city, or it could be implemented in a smartphone, operated by individuals wishing to prove their location. It is also worth noting that T-PoP is not a highly resource intensive algorithm to execute, and thus the hardware constraints are not a critical concern. It is comprised of two algorithms: Verification and Checks. These are further defined in 3 and 2 accordingly. Verification is a (reverse) breadth-first-search algorithm, meaning its runtime in the worst case scenario is $O(|V|+|E|)$, where $V$ is the number of nodes and $E$ the number of edges in the tree data structure [24]. The Verification algorithm calls the Checks algorithm at each node, and the Checks algorithm has a time complexity of $O(|V|)$. So T-PoP’s time complexity is $O(|V|^{2}+|E||V|)$.
@@ VII-A Agent States
In practical terms: an honest agent will always claim to be in their real position. If $P(\alpha_{i}=\overline{h})$, finding $P(\gamma_{i}=t)$ is slightly more convoluted. This is because when a dishonest agent is sampled in a tree, one cannot, in theory, distinguish if its claimed position is its real position or not. Indeed, that is the purpose of T-PoP: to determine whether an agent is claiming to be in their real position.
@@ VII-B2 Picking a child, given a parent
$x(\textbf{c})=s_{v}=\{\alpha_{\textbf{c}},\beta_{\textbf{c}},\gamma_{\textbf{c}}\}$
$x(\textbf{p})=s_{u}=\{\alpha_{\textbf{p}},\beta_{\textbf{p}},\gamma_{\textbf{p}}\}$
To compute $m_{u,v}$ it suffices to evaluate:
@@
Note that only the coercion attribute of the parent affects the result of equations 18. We include all three attributes to ease exposition, to show all possible state combinations.
@@ VII-C Condition 2: The tree has sufficient nodes
This means, that for TPoP with a certain $\theta$ the probability of satisfying condition 2, $P(\text{Criteria $2$ $=$ True}|s_{u},\theta)$, for all $s_{u}$, can be expressed as
$\prod_{i=0}^{h-1}\prod_{j=1}^{n_{i}}(1-\sum_{k=0}^{n_{i+1}-1}e^{-\lambda}\frac{\lambda^{k}}{k!}),$ (26)
which is simply the product of the probabilities that each parent in the tree will have enough neighbours.
@@ VII-D Condition 3: all the nodes in the tree are unique
  • TPoP with $\theta_{1}=(t,h=1,w_{1}=6)$
  • TPoP with with $\theta_{2}=(t,h=2,w_{1}=2,w_{2}=2)$.
For brevity, we denote $P(\text{Criteria $3$ $=$ True}|s_{u},\theta)$ as $p(\theta)$ (as the probability of condition 3 being true does not depend on the state of the prover). In the remainder of this section we make the assumption that each agent selects witnesses among its neighbours without replacement. Although it may be an unrealistic assumption (as in practice each prover would select distinct witnesses and they would not deliberately try to hinder their own proof), it allows us to provide a lower bound on the probability of selecting a tree made of unique nodes.
@@
For the case of parameter $\theta_{1}$, $p(\theta_{1})$ is straightforward to compute and it is equal to $\kappa_{0,6}$.
For the tree with parameter $\theta_{2}$, finding the probability that all nodes are unique is more complex. Let us consider a tree started by a prover $a_{1}$ at the root, tree height of $h=2$, and branching factor of $w=2$, i.e., each parent choosing two of its neighbours at random (with replacement) as its children (i.e., the nodes at level 1 will not choose themselves).
@@
Finally, inserting equation (31) into equation (28), we have the theoretical approximation for the probability of the tree being comprised of strictly distinct agents, i.e., $p(\theta_{2})=P(L_{0},L_{1},L_{2})=\kappa_{0,2}\cdot P(L2|L_{0},L_{1})$. This value of $P(L_{0},L_{1},L_{2})$, given a constant range of sight $r$, is only dependent on the density of agents $N/l^{2}$ and can be calculated numerically.
@@
As a final remark we note that it is very difficult to to obtain a general expression to compute the probability of finding distinct agents in a tree with for any $\theta$. It is however quite straightforward to approximate $p(\theta)$, for a given $\langle k\rangle$ and a given $r$, by making use of a large number of Monte-Carlo simulations.
@@ IX Agent Based Simulations
We construct an agent-based simulator, coded in Python, which is available in our GitHub Repository. In it, we create an environment of a given size, and instantiate a fixed number of agents with a fixed value for their range of sight. Agents are assigned an honesty, $\alpha_{i}$, and a coercion, $\beta_{i}$, attribute independently and at random, following equations 10a and 10b. We repeat the simulations for values of $p_{h}$ and $p_{c}$ belonging to the interval $[0,1]$. All agents are given a real position, within the bounds of the defined environment, and dishonest agents also have a claimed position that is different to their real position. These positions are uniformly and randomly distributed across the environment. We first simulate a low-density scenario, as shown in Figures 17, 16, 19 and 18. We initialise an environment of size 10 by 10 arbitrary square units, and instantiate 350 agents, such that each has a range of sight of 1 unit. Thus, agents have an average of 3.5 agents within their field of view. Each agent constructs a tree, and then runs the T-PoP algorithms to obtain a proof-of-position. These simulations are carried out across the ranging values of $p_{h}$ and $p_{c}$, starting from 0 to 100% in intervals of 10%. At each combination of $p_{h}$ and $p_{c}$, 5 Monte Carlo simulations are run for every agent. In total, that is 211,750 T-PoP simulations [^4] There are 121 possible combinations of $p_{h}$ and $p_{c}$ when varying the values at intervals of 10%, 350 agents and 5 simulations per possible combination of $p_{h}$ and $p_{c}$ yields $5*121*350$.. We test the performance of T-PoP for the following operating conditions:
$\theta_{1}=(t=1,w_{1}=2,w_{2}=2,h=2)$
$\theta_{2}=(t=1,w_{1}=6,h=1)$
$\theta_{3}=(t=0.4,w_{1}=2,w_{2}=2,h=2)$
$\theta_{4}=(t=0.4,w_{1}=6,h=1)$
In order to assess the performance of TPoP for each scenario we use an empirical approximation of the True Negatives (Security or TN) and of the True Positives (reliability ot TP). For completeness we also consider the False Negatives (FN) and the False Positives (FP). Each classification instance is described as below:
@@
Notice that we consider recall over precision because we are particularly interested in the sensitivity of the algorithm. Namely, it is of utmost importance that the number of True Negatives is high, as this value represents the security of the algorithm (i.e., how easy it is for a dishonest agent to be classified as honest).
@@
Remark: we note that the results of the simulations, illustrated in images 16-23, present artifacts for $p_{h}$ belonging to $[0,0.1]$ and $[0.9,0.1]$. This is due to the fact that in those ranges there are either no honest agents and no honest agents (due to the resolution of $10\%$). We note however that the results, aside from these artifacts, are consistent with the theoretical results presented in Figures 8 and 9.
@@ X Performance of TPoP
From the results obtained, it can be observed that the performance of TPoP does not vary greatly when using parameter $h=1,w_{1}=6$ versus $h=2,w_{1}=2,w_{2}=2$. The former slightly out-performs the latter in both thresholds tested. This is consistent with the results shown in Figures 10. The main difference in performance is observed when varying the threshold parameter, $t$, and when varying the density. Increasing density increases the number of True Positives, given that honest agents will not be classified as dishonest for not having sufficient nodes in their tree. Decreasing the threshold improves the rate of True Positives (Reliability) of T-PoP in very low density scenarios, but this comes at the expense of Security. A lower threshold makes it easier for dishonest agents to submit a valid tree. As such, users may tune the threshold parameters to obtain a desired level of Security or Reliability, depending on the expected density of their system.